# Rho-1: Not All Tokens Are What You Need

### v1

@@ Abstract
Previous language model pre-training methods have uniformly applied a next-token prediction loss to all training tokens. Challenging this norm, we posit that “Not all tokens in a corpus are equally important for language model training”. Our initial analysis delves into token-level training dynamics of language model, revealing distinct loss patterns for different tokens. Leveraging these insights, we introduce a new language model called Rho-1. Unlike traditional LMs that learn to predict every next token in a corpus, Rho-1 employs Selective Language Modeling (SLM), which selectively trains on useful tokens that aligned with the desired distribution. This approach involves scoring pretraining tokens using a reference model, and then training the language model with a focused loss on tokens with higher excess loss. When continual pretraining on 15B OpenWebMath corpus, Rho-1 yields an absolute improvement in few-shot accuracy of up to 30% in 9 math tasks. After fine-tuning, Rho-1-1B and 7B achieved state-of-the-art results of 40.6% and 51.8% on MATH dataset, respectively — matching DeepSeekMath with only 3% of the pretraining tokens. Furthermore, when pretraining on 80B general tokens, Rho-1 achieves 6.8% average enhancement across 15 diverse tasks, increasing both efficiency and performance of the language model pre-training.
@@ 1 Introduction
To explore how language models learn at the token level, we initially examined training dynamics, particularly how the token-level loss evolves during usual pretraining. In §2.1, we evaluated the model’s token perplexity at different checkpoints and categorized tokens into different types. Our findings reveal that significant loss reduction is limited to a select group of tokens during training. Many tokens are “easy tokens” that are already learned, and some are “hard tokens” that exhibit variable losses and resist convergence. These tokens can lead to numerous ineffective gradient updates.
@@
We show through comprehensive experiments that SLM significantly enhances token efficiency during pretraining and improves performance on downstream tasks. Furthermore, our findings indicate that SLM effectively identifies tokens relevant to the target distribution, resulting in improved perplexity scores on benchmarks for models trained with the selected tokens. §3.2 shows the effectiveness of SLM on math continual pretraining: both 1B and 7B Rho-1 outperform CLM-trained baselines by over 16% on the GSM8k and MATH datasets. SLM reaches baseline accuracy up to 10x faster, as shown in Figure 1. Remarkably, Rho-1-7B matches the state-of-the-art performance of DeepSeekMath-7B using only 15B tokens, compared to the 500B tokens required by DeepSeekMath. Upon fine-tuning, Rho-1-1B and 7B achieve 40.6% and 51.8% on MATH, respectively. Notably, Rho-1-1B is the first 1B LM to exceed 40% accuracy, nearing the early GPT-4’s CoT performance of 42.5%. §3.3 confirms the efficacy of SLM in general pretraining: Training Tinyllama-1B on 80B tokens with SLM improves 6.8% on average across 15 benchmarks, with gains over 10% in code and math tasks.
@@ 2.1 Not All Tokens Are Equal: Training Dynamics of Token Loss
Our investigation begins with a critical look at how individual tokens’ losses evolve during standard pre-training. We continue pre-training Tinyllama-1B with 15B tokens from OpenWebMath, saving checkpoints after every 1B tokens. We then evaluate token-level loss at these intervals using the validation set of approximately 320,000 tokens. Figure 3(a) reveals a striking pattern: tokens fall into four categories based on their loss trajectory—persistent high loss (H→H), increasing loss (L→H), decreasing loss (H→L), and consistent low loss (L→L). For further details on these categories, see §B.1. Our analysis uncovers that a mere 26% of tokens show a notable loss reduction (H→L), while the majority (51%) remain in the L→L category, indicating they have already been learned. Interestingly, 11% of the tokens are persistently challenging (H→H), likely due to high aleatoric uncertainty (Hüllermeier and Waegeman, 2021). Additionally, 12% of tokens experience an unexpected loss increase (L→H) during training.
Our second observation is that a significant number of token losses exhibit persistent fluctuations, and resist convergence. The loss of many L→L and H→H tokens, as depicted in Figure 3 (b) and (c), show high variance during training. In §B.2, we visualize and analyze the content of these tokens and find that many of them are noisy, which is consistent with our hypothesis.
@@ 2.2 Selective Language Modeling
We begin by curating a high-quality dataset that reflects the desired data distribution. We train a reference model (RM) using standard cross-entropy loss on the curated data. The resulting RM is then used to assess the token loss within a larger pretraining corpus. We compute the reference loss ($\mathcal{L}_{\text{ref}}$) of a token $x_{i}$ based on the probability that the RM assigns to this token. The calculation is formalized as follows:
$\mathcal{L}_{\text{ref}}(x_{i})=-\log P(x_{i}|x{<i})$ (1)
By evaluating $\mathcal{L}_{\text{ref}}$ for each token, we establish the reference loss for selective pretraining, allowing us to focus on the most influential tokens in language modeling.
Figure 4: The pipeline of Selective Language Modeling (SLM). SLM optimizes language model performance by concentrating on valuable, clean tokens during pre-training. It involves three steps: (Step 1) Initially, train a reference model on high-quality data. (Step 2) Then, score each token’s loss in a corpus using the reference model. (Step 3) Finally, train the language model selectively on tokens that show higher excess loss compared to the reference loss.
@@
$\mathcal{L}_{\Delta}(x_{i})=\mathcal{L}_{\theta}(x_{i})-\mathcal{L}_{\text{ref}}(x_{i})$ (3)
@@
$I_{k\%}(x_{i})=\begin{cases}1&\text{if }x_{i}\text{ is in the top }k\%\text{ of }\mathcal{L}_{\Delta}\\ 0&\text{otherwise}\end{cases}$ (5)
This ensures that the loss is applied only to the tokens that are deemed most beneficial for the language model to learn from. In practice, token selection can be implemented by ranking the tokens in a batch according to their excess loss and using only the top $k\%$ of tokens for training. This process eliminates the loss for undesired tokens without incurring additional costs during pretraining, making our approach both efficient and easily integrated.
@@ 3.1 Experimental Setup
Table 1: Few-shot CoT reasoning results of math pretraining. All models are tested with few-shot prompting. Previous best results are highlighted in blue, while our best results are in purple. $*{}^{*}$nly unique math-related tokens are calculated. For Rho-1, we calculate only the selected tokens that are used for training. ${}^{\dagger}$e use OpenAI’s MATH subset (Lightman et al., 2023) for evaluation, since some original test samples have been used in public training sets such as PRM800k. ${}^{\ddagger}$he SAT only has 32 four-choice problems, so we average our results over the last three checkpoints, if available.
@@
GSM8K ${}^{\dagger}$ SVAMP ASDiv MAWPS TAB MQA
@@ 1-2B Base Models
For math pretraining, we continue pretraining on the Tinyllama-1.1B model (Zhang et al., 2024) and the Mistral-7B model (Jiang et al., 2023) with learning rates of 8e-5 and 2e-5, respectively. For general domain, we set the learning rate for Tinyllama-1.1B model to 1e-4. The batch size is uniformly set to 1M tokens for both domains. Regarding the token selection ratio, we use 60% for the Tinyllama-1.1B model and 70% for the Mistral-7B model.
@@
We use models that have been continually pretrained (Tinyllama-CT and Mistral-CT) through regular causal language modeling as baselines. Moreover, we compare Rho-1 with well-known and top-performing baselines, including Gemma (Team et al., 2024), Qwen1.5 (Bai et al., 2023), Phi-1.5 (Li et al., 2023b), DeepSeekLLM (DeepSeek-AI, 2024), DeepSeekMath (Shao et al., 2024a), CodeLlama (Roziere et al., 2023), Mistral (Jiang et al., 2023), Minerva (Lewkowycz et al., 2022), Tinyllama (Zhang et al., 2024), LLemma (Azerbayev et al., 2023), and InternLM2-Math (Ying et al., 2024). For fine-tuning results, we also compare with previous best models MAmmoTH(Yue et al., 2024) and ToRA(Gou et al., 2024).
@@
To comprehensively evaluate pretrained models, we compare their few-shot capabilities and fine-tuning performance across a variety of tasks. We adopt the lm-eval-harness [^1] https://github.com/EleutherAI/lm-evaluation-harness for general tasks, and math-eval-harness [^2] https://github.com/ZubinGou/math-evaluation-harness for math tasks. We use vllm (v0.3.2) (Kwon et al., 2023) to speed up inference. Further details on our evaluation can be found in Appendix C.
@@ 3.2 Math Pre-training Results
We evalute base models prompting with few-shot chain-of-thought (CoT) (Wei et al., 2022a) examples following previous works (Lewkowycz et al., 2022; Azerbayev et al., 2023; Shao et al., 2024b). As results shown in Table 1, in comparison to continue pretraining directly, Rho-1-Math has achieved the average few-shot accuracy improvement of 16.5% on 1B models and 10.4% on 7B models. Furthermore, after training for multiple epochs on OpenWebMath, we find that Rho-1 could further increase the average few-shot accuracy to 40.9%. Compared to DeepSeekMath-7B, which pretrained on 500 billion math-related tokens, Rho-1-7B pretrained on only 15 billion tokens (selecting 10.5 billion tokens) achieved comparable results, demonstrating the efficiency of our approach.
@@ 3.3 General Pre-training Results
Figure 6: The dynamics of pretraining loss and downstream loss. (a) and (c) represent the loss of tokens selected/unselected by SLM during pretraining in both SLM and CLM methods, while (b) represents the loss of the SLM and CLM methods on downstream corpora. We tested the above results through the process of pretraining with a total of 4 billion tokens.
3.4 Analysis
@@
We use the reference model to filter tokens and explore the changes in validation loss after training on all/selected tokens, while observing their relationship with downstream loss. As shown in Figure 6, we pretrain approximately 4B tokens and displayed the variation curves of loss on different pretraining method and validation sets during the pretraining process. We can observe that on the tokens selected by the reference model, the decrease in average loss of the Rho-1 is more significant compared to regular pretraining. On the contrary, on the unselected tokens, the decrease in average loss of the regular pretraining is more significant. If we associate fig (a), fig(b) with fig(c), it is not difficult to find that the model trained on the selected tokens has a more significant decrease in downstream loss, while ordinary pretraining, although reducing the average loss of all tokens during the training phase, is difficult to have a significant decrease in downstream loss. Therefore, we expect that selecting tokens for pretraining is more efficient.
@@
Moreover, We relate the selected tokens’ loss to its downstream task performance via a power law in Figure 7, which is similar to a concurrent study (Gadre et al., 2024). Observing the curve fitted from the data points in the graph, the average loss of tokens selected by the SLM shows a positive correlation with the performance of downstream tasks, whereas the average loss of tokens not selected exhibits a negative correlation with downstream task performance. Therefore, it is not necessary for the all tokens’ loss to decrease to benefit the model’s ultimate performance. See Appendix D for more details.
[Uncaptioned image] Figure 8: The PPL of tokens selected by different checkpoint. We test the PPL of the tokens selected at 2B, 5B, 8B, 11B, and 14B.
[Uncaptioned image] Figure 9: Effect of token select ratio. We train 1B LM with SLM objective on 5B tokens.
@@
We aim to analyze the tokens selected by the SLM method in pretraining to further explore its working mechanism. To this end, we visualize the token selection process during the training of Rho-1 using the OpenWebMath. In §E.1, we have highlighted in blue the tokens that were retained during actual pretraining. We observe that the majority of tokens chosen by the SLM method are closely related to mathematics, effectively training the model on the parts of the original corpus that are pertinent to mathematical content.
@@
Weak-to-Strong Generization
Table 3: Weak-to-Strong generization result on math benchmark.
Model GSM8K ${}^{\dagger}$ SVAMP ASDiv MAWPS TAB MQA
MMLU
STEM
${}^{\ddagger}$ AVG
Llama-2-7B-CT 28.4 13.6 50.3 62.8 79.5 37.6 34.1 41.6 43.5 43.5
Llama-2-7B-CT w/ 1B RM 29.8 16.0 55.5 63.7 80.4 37.9 34.3 38.2 43.8 44.4
Apart from the main experiments where we use the same base model for the reference and continual pretraining, we also investigate if a smaller reference model can effectively guide the pretraining of a larger model. We use Tinyllma-1.1B as reference model and continual pretraining Llama-2-7B on math. Results presented in Table 3 indicate that, despite the considerable gap between the small and large models (Li et al., 2023c), employing the small reference model to token selection can still yield benefits to the pre-training of the larger model. If reference and training models have different vocabularies, one can consider performing token alignment (Wan et al., 2024; Fu et al., 2023), which we leave for future work.
4 Related Works
Pretraining Data Optimization

### v2

@@ Abstract
Previous language model pre-training methods have uniformly applied a next-token prediction loss to all training tokens. Challenging this norm, we posit that “Not all tokens in a corpus are equally important for language model training”. Our initial analysis examines token-level training dynamics of language model, revealing distinct loss patterns for different tokens. Leveraging these insights, we introduce a new language model called Rho-1. Unlike traditional LMs that learn to predict every next token in a corpus, Rho-1 employs Selective Language Modeling (SLM), which selectively trains on useful tokens that aligned with the desired distribution. This approach involves scoring pretraining tokens using a reference model, and then training the language model with a focused loss on tokens with higher scores. When continual pretraining on 15B OpenWebMath corpus, Rho-1 yields an absolute improvement in few-shot accuracy of up to 30% in 9 math tasks. After fine-tuning, Rho-1-1B and 7B achieved state-of-the-art results of 40.6% and 51.8% on MATH dataset, respectively — matching DeepSeekMath with only 3% of the pretraining tokens. Furthermore, when pretraining on 80B general tokens, Rho-1 achieves 6.8% average enhancement across 15 diverse tasks, increasing both efficiency and performance of the language model pre-training.
@@ 1 Introduction
To explore how language models learn at the token level, we initially examined training dynamics, particularly how the token-level loss evolves during usual pretraining. In §2.1, we evaluated the model’s token perplexity at different checkpoints and categorized tokens into different types. Our findings reveal that significant loss reduction is limited to a select group of tokens. Many tokens are “easy tokens” that are already learned, and some are “hard tokens” that exhibit variable losses and resist convergence. These tokens can lead to numerous ineffective gradient updates.
@@
We show through comprehensive experiments that SLM significantly enhances token efficiency during pretraining and improves performance on downstream tasks. Furthermore, our findings indicate that SLM effectively identifies tokens relevant to the target distribution, resulting in improved perplexity scores on benchmarks for models trained with the selected tokens. §3.2 shows the effectiveness of SLM on math continual pretraining: both 1B and 7B Rho-1 outperform CLM-trained baselines by over 16% on the GSM8k and MATH datasets. SLM reaches baseline accuracy up to 10x faster, as shown in Figure 1. Remarkably, Rho-1-7B matches the state-of-the-art performance of DeepSeekMath-7B using only 15B tokens, compared to the 500B tokens required by DeepSeekMath. Upon fine-tuning, Rho-1-1B and 7B achieve 40.6% and 51.8% on MATH, respectively. Notably, Rho-1-1B is the first 1B LM to exceed 40% accuracy, nearing the early GPT-4’s CoT performance of 42.5%. §3.3 confirms the efficacy of SLM in general pretraining: Training Tinyllama-1B on 80B tokens with SLM improves 6.8% on average across 15 benchmarks, with gains over 10% in code and math tasks. In §3.4, we demonstrate that in settings without high-quality reference data, we can use SLM for self-referencing, leading to an average improvement of up to 3.3% in downstream tasks.
@@ 2.1 Not All Tokens Are Equal: Training Dynamics of Token Loss
Our investigation begins with a critical look at how individual tokens’ losses evolve during standard pre-training. We continue pre-training Tinyllama-1B with 15B tokens from OpenWebMath, saving checkpoints after every 1B tokens. We then evaluate token-level loss at these intervals using the validation set of approximately 320,000 tokens. Figure 3(a) reveals a striking pattern: tokens fall into four categories based on their loss trajectory—persistent high loss (H→H), increasing loss (L→H), decreasing loss (H→L), and consistent low loss (L→L). For further details on these categories, see §D.1. Our analysis uncovers that a mere 26% of tokens show a notable loss reduction (H→L), while the majority (51%) remain in the L→L category, indicating they have already been learned. Interestingly, 11% of the tokens are persistently challenging (H→H), likely due to high aleatoric uncertainty (Hüllermeier and Waegeman, 2021). Additionally, 12% of tokens experience an unexpected loss increase (L→H) during training.
Our second observation is that a significant number of token losses exhibit persistent fluctuations, and resist convergence. The loss of many L→L and H→H tokens, as depicted in Figure 3 (b) and (c), show high variance during training. In §D.2, we visualize and analyze the content of these tokens and find that many of them are noisy, which is consistent with our hypothesis.
@@ 2.2 Selective Language Modeling
We begin by curating a high-quality dataset that reflects the desired data distribution. We train a reference model (RM) using standard cross-entropy loss on the curated data. The resulting RM is then used to assess the token loss within a larger pretraining corpus. We compute the reference loss ($\mathcal{L}_{\text{RM}}$) of a token $x_{i}$ based on the probability that the RM assigns to this token. The calculation is formalized as follows:
$\mathcal{L}_{\text{RM}}(x_{i})=-\log P(x_{i}|x_{<i})$ (1)
By evaluating $\mathcal{L}_{\text{RM}}$ for each token, we establish the reference loss for selective pretraining, allowing us to focus on the most influential tokens in language modeling.
Figure 4: The pipeline of Selective Language Modeling (SLM). SLM optimizes language model performance by concentrating on valuable, clean tokens during pre-training. It involves three steps: (Step 1) Initially, train a reference model on high-quality data. (Step 2) Then, score each token’s loss in a corpus using the reference model. (Step 3) Finally, selectively train the language model on tokens that have higher scores.
@@
$\mathcal{L}_{\Delta}(x_{i})=\mathcal{L}_{\theta}(x_{i})-\mathcal{L}_{\text{RM}}(x_{i})$ (3)
@@
$I_{k\%}(x_{i})=\begin{cases}1&\text{if }x_{i}\text{ ranks in the top }k\%\text{ by }S(x_{i})\\ 0&\text{otherwise}\end{cases}$ (5)
By default, we use $\mathcal{L}_{\Delta}$ as the score function $S$. This ensures that the loss is applied only to the tokens that are deemed most beneficial for the language model to learn from. In practice, token selection can be implemented by ranking the tokens in a batch according to their excess loss and using only the top $k\%$ of tokens for training. This process eliminates the loss for undesired tokens without incurring additional costs during pretraining, making our approach both efficient and easily integrated.
@@ 3.1 Experimental Setup
Table 1: Few-shot CoT reasoning results of math pretraining. All models are tested with few-shot prompting. Previous best results are highlighted in blue, while our best results are in purple. Only unique math-related tokens are calculated. For Rho-1, we calculate only the selected tokens that are used for training. We use OpenAI’s MATH subset (Lightman et al., 2023) for evaluation, since some original test samples have been used in public training sets such as PRM800k. The SAT only has 32 four-choice problems, so we average our results over the last three checkpoints, if available.
@@
GSM8K MATH SVAMP ASDiv MAWPS TAB MQA
@@ 1-2B Base Models
For math pretraining, we continue pretraining on the Tinyllama-1.1B model (Zhang et al., 2024) and the Mistral-7B model (Jiang et al., 2023) with learning rates of 8e-5 and 2e-5, respectively. For the 1.1B model, we conducted our training on 32 × H100 80G GPUs. This configuration allowed us to train approximately 15 billion tokens in around 3.5 hours and 50 billion tokens in about 12 hours. In the case of the 7B model, training the same 15 billion tokens took approximately 18 hours under similar hardware conditions. For general domain, we set the learning rate for Tinyllama-1.1B model to 1e-4 and train 80B tokens under the same hardware conditions, which takes approximately 19 hours. The batch size is uniformly set to 1M tokens for both domains. Regarding the token selection ratio, we use 60% for the Tinyllama-1.1B model and 70% for the Mistral-7B model.
@@
We use models that have been continually pretrained (Tinyllama-CT and Mistral-CT) through regular causal language modeling as baselines. Moreover, we compare Rho-1 with well-known and top-performing baselines, including Gemma (Team et al., 2024), Qwen1.5 (Bai et al., 2023), Phi-1.5 (Li et al., 2023b), DeepSeekLLM (DeepSeek-AI, 2024), DeepSeekMath (Shao et al., 2024), CodeLlama (Roziere et al., 2023), Mistral (Jiang et al., 2023), Minerva (Lewkowycz et al., 2022), Tinyllama (Zhang et al., 2024), LLemma (Azerbayev et al., 2023), and InternLM2-Math (Ying et al., 2024). For fine-tuning results, we also compare with previous best models MAmmoTH(Yue et al., 2024) and ToRA(Gou et al., 2024).
@@
To comprehensively evaluate pretrained models, we compare their few-shot capabilities and fine-tuning performance across a variety of tasks. We adopt the lm-eval-harness [^1] https://github.com/EleutherAI/lm-evaluation-harness (Gao et al., 2023) for general tasks, and develop math evaluation suite [^2] https://github.com/ZubinGou/math-evaluation-harness for math tasks. We use vllm (v0.3.2) (Kwon et al., 2023) to speed up inference. Further details on our evaluation can be found in Appendix E.
@@ 3.2 Math Pre-training Results
We evalute base models prompting with few-shot chain-of-thought (CoT) (Wei et al., 2022a) examples following previous works (Lewkowycz et al., 2022; Azerbayev et al., 2023; Shao et al., 2024). As results shown in Table 1, in comparison to continue pretraining directly, Rho-1-Math has achieved the average few-shot accuracy improvement of 16.5% on 1B models and 10.4% on 7B models. Furthermore, after training for multiple epochs on OpenWebMath, we find that Rho-1 could further increase the average few-shot accuracy to 40.9%. Compared to DeepSeekMath-7B, which pretrained on 500 billion math-related tokens, Rho-1-7B pretrained on only 15 billion tokens (selecting 10.5 billion tokens) achieved comparable results, demonstrating the efficiency of our approach.
@@ 3.3 General Pre-training Results
3.4 Self-Reference Results
Table 3: Self-Reference results. We use OpenWebMath (OWM) to train the reference model.
Model
Score
Function
Data
Uniq.
Toks
Train
Toks
GSM8K MATH SVAMP ASDiv MAWPS MQA AVG
Tinyllama-CT (RM) - OWM 14B 15B 6.3 2.6 21.7 36.7 47.7 13.9 21.5
Tinyllama-SLM $\mathcal{L}_{\text{RM}}$ OWM 14B 10.5B 6.7 4.6 23.3 40.0 54.5 14.3 23.9
Tinyllama-SLM $\mathcal{H}_{\text{RM}}$ OWM 14B 10.5B 7.0 4.8 23.0 39.3 50.5 13.5 23.0
Tinyllama-SLM $\mathcal{L}_{\text{RM}}\cap\mathcal{H}_{\text{RM}}$ OWM 14B 9B 7.1 5.0 23.5 41.2 53.8 18.0 24.8
Tinyllama-CT - PPile 55B 52B 8.0 6.6 23.8 41.0 54.7 14.2 24.7
Tinyllama-SLM $\mathcal{L}_{\text{RM}}\cap\mathcal{H}_{\text{RM}}$ PPile 55B 36B 8.6 8.4 24.4 43.6 57.9 16.1 26.5
In this section, we demonstrate that SLM can enhance the effectiveness of model pre-training using only pre-training corpora, without the need for additional high-quality data. Specifically, we initially trained the reference model on the OpenWebMath (OWM) corpus, a subset of Proof-Pile-2 (PPile). We evaluated OWM and PPile using the trained reference model and selected tokens for training. In this scenario, we assume the absence of downstream task-related data, a common situation in real-world applications. We hypothesize that the key factor is not scoring the desired distribution but filtering out noisy tokens. Therefore, we employed two different scoring functions based on the reference model loss, $\mathcal{L}_{\text{RM}}$, and the information entropy of the next token, $\mathcal{H}_{\text{RM}}$, which measures the uncertainty of the next token. Details are provided in Appendix H.
The experimental results, as shown in Table 3, indicate that using only the OWM-trained reference model can effectively guide the model in pre-training on the same corpus, improving average downstream performance by +2.4%. Using only the information entropy as the score function brought about a similar improvement. Additionally, we considered training on the intersection of tokens selected by the two scoring functions and found better performance, with a 40% reduction in tokens and +3.3% performance. Furthermore, training the SLM on the PPile, despite only using the OWM subset to train the reference model, still achieved a 1.8% improvement with 30% fewer tokens used. For more details, please refer to Appendix H.
Figure 6: The dynamics of pretraining loss and downstream loss. (a) and (c) represent the loss of tokens selected/unselected by SLM during pretraining in both SLM and CLM methods, while (b) represents the loss of the SLM and CLM methods on MetaMath (Yu et al., 2024). We tested the above results through the process of pretraining with a total of 4 billion tokens.
3.5 Ablation Study and Analysis
@@
We utilized the reference model to filter tokens and assess their impact on validation and downstream losses after training. As depicted in Figure 6, we pretrained on 4B tokens and tracked loss variations across methods and validation sets. The Rho-1 showed greater loss reduction on selected tokens than regular pretraining. Cross-referencing figures (a), (b), and (c) reveals that selected-token pretraining substantially lowers downstream loss, while traditional pretraining’s effect on downstream loss is less pronounced despite initial loss reductions. Therefore, we expect that selecting tokens for pretraining is more efficient.
@@
In Figure 7, we demonstrate that the loss of selected tokens correlates with downstream task performance, following a power law similar to recent findings (Gadre et al., 2024). Our analysis shows that tokens selected by SLM positively impact performance, while those not selected have a negative impact. Thus, reducing loss across all tokens is not imperative for improved model performance. Refer to Appendix F for further details.
Figure 8: The PPL of tokens selected by different checkpoint. We test the PPL of the tokens selected at 2B, 5B, 8B, 11B, and 14B.
Figure 9: Effect of token select ratio. We train 1B LM with SLM objective on 5B tokens.
@@
We aim to analyze the tokens selected by the SLM method in pretraining to further explore its working mechanism. To this end, we visualize the token selection process during the training of Rho-1 using the OpenWebMath. In §G.1, we have highlighted in blue the tokens that were retained during actual pretraining. We observe that the majority of tokens chosen by the SLM method are closely related to mathematics, effectively training the model on the parts of the original corpus that are pertinent to mathematical content.
@@
4 Conclusion
In this paper, we propose using Selective Language Modeling(SLM) to train Rho-1, which select more suitable tokens for current pretraining stage. We conducted the detailed analysis of the loss of tokens during the pretraining process and found that not all tokens are equal during pretraining. Our experiments and analysis in the fields of mathematics and general have demonstrated the effectiveness of the SLM method, emphasizing the importance of token level in the LLM pretraining process. In the future, how to improve pretraining of LLMs from the perspective of token level worthy of in-depth research.
Appendix
[…]
Appendix A Author Contributions
Appendix B Related Works
B.1 Pretraining Data Optimization